The following is the equation which connects u, p and v for several gases

u = a + bpv

where a and b are constants. Prove that for a reversible adiabatic process,

pv^γ = constant, where γ = b +1/b


1
Expert's answer
2021-08-25T09:12:07-0400

For a reversible adiabatic process

"Q=0"

For a process of unit mass system

"dq=du+pdv"

or

"0=du+pdv"

We have

"u=a+bpv"

Differentiating both sides,

"du=bvdp+bpdv \\

0=bvdp+bpdv+pdv \\

p(1+b)dv + bvdp=0"

Multiplying both sides by "\frac{1}{bvp}" , we get

"\frac{1+b}{b} \times \frac{dv}{v} + \frac{dp}{p}=0"

Integrating both sides

"(\frac{1+b}{b})ln(v) +ln(p) = ln(C)"

Taking antilog

"pv^{(\frac{1+b}{b})}=C"

Using

"γ= \frac{1+b}{b}"

then

"pv^\u03b3=C"

Proved.